Review Part A Find an expression for the speed of the ice at the bottom. Express your answer in terms of some or all of the variables h, 8, and free-fall acceleration g. A very slippery block of ice slides down a smooth ramp tilted at angle 6. The ice is released from rest at vertical height h above the bottom of the ramp. Submit Provide Feedback
Answer #1

Given,

angle = theta ; height = h

from conservation of energy

PE = Ke

m g h = 1/2 m v^2

v = sqrt (2 g h)

Its independent of angle ;

One more approach to the solution:

the acceleration of the block will be:

a = g sin(theta)

distance moved is s, using trigonometry

s = h/sin(theta)

We know from eqn of motion

v^2 = u^2 + 2 a s

v^2 = 0^2 + 2 x g sin(theta) x h/sin(theta)

v = sqrt (2 g h)